# Existence of solution to a system of PDEs modeling the crystal growth inside lithium batteries

### v2

@@ 5.1. Finite element approximation
$V_{h}=\{v\in\operatorname{C}^{0}(\bar{\Omega});\leavevmode\nobreak\ v|_{\mathcal{T}}\in P_{2}(\mathcal{T}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \mathcal{T}\in\mathscr{T}_{h}\}\subset\operatorname{H}^{1}(\Omega),$
$V_{0,h}=\{v\in\operatorname{C}^{0}(\bar{\Omega});\leavevmode\nobreak\ v=0\text{ on }\Sigma\subseteq\partial\Omega;\leavevmode\nobreak\ v|_{\mathcal{T}}\in P_{2}(\mathcal{T}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \mathcal{T}\in\mathscr{T}_{h}\}\subset\operatorname{H}^{1}_{0|\Sigma}(\Omega),$
where $\Sigma=\Gamma_{1}\cup\Gamma_{2}$ and $P_{2}(\mathcal{T})$ denotes the set of polynomials of maximum degree $2$ on $\mathcal{T}$.
@@
$\epsilon^{2}(d_{\tau}u^{k}_{h},\chi_{h})+(\boldsymbol{{A}}(\nabla u^{k-1}_{h})\nabla u^{k}_{h},\nabla\chi_{h})+\gamma(g^{\prime}(u^{k-1}_{h}),\chi_{h})=(\tilde{m}(c^{k-1}_{h})h^{\prime}(u^{k-1}_{h}),\chi_{h}),$ (5.1a)
@@
where $\tilde{m}(c)=C_{1}(e^{-C_{2}}-ce^{C_{2}})$. The physical parameters are defined, together with their values in the computations, in Table 1. We note that (5.1a) – (5.1c) are linear in $u^{k}$, $\phi^{k}$ and $c^{k}$ respectively.
@@ 5.2. Numerical examples
$\hskip 85.35826pt-(\beta u_{h}^{k}(1-u^{k}_{h})X,\chi_{h}),\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ \chi_{h}\in V_{h},$ (5.7)
@@
Electrochemical reaction rate coefficients $C_{1}$ $C_{2}$ $1/15$ $1.9345$

### v3

@@ 5.1. Finite element approximation
$V_{h}=\{v\in\operatorname{C}^{0}(\bar{\Omega});\penalty 10000\ v|_{\mathcal{T}}\in P_{2}(\mathcal{T}),\penalty 10000\ \forall\penalty 10000\ \mathcal{T}\in\mathscr{T}_{h}\}\subset\operatorname{H}^{1}(\Omega),$
$V_{0,h}=\{v\in\operatorname{C}^{0}(\bar{\Omega});\penalty 10000\ v=0\text{ on }\Sigma\subseteq\partial\Omega;\penalty 10000\ v|_{\mathcal{T}}\in P_{2}(\mathcal{T}),\penalty 10000\ \forall\penalty 10000\ \mathcal{T}\in\mathscr{T}_{h}\}\subset\operatorname{H}^{1}_{0|\Sigma}(\Omega),$
where $\Sigma=\Gamma_{1}\cup\Gamma_{2}$ and $P_{2}(\mathcal{T})$ denotes the set of polynomials of maximum degree $2$ on $\mathcal{T}$. $P_{1}(\mathcal{T})$ would be an alternative choice, but $P_{2}(\mathcal{T})$ elements lead to smoother visual results at interfacial layers.
@@
$\begin{split}\epsilon^{2}(d_{\tau}u^{k}_{h},\chi_{h})+(\boldsymbol{{A}}(\nabla u^{k-1}_{h})\nabla u^{k}_{h},\nabla\chi_{h})+\gamma(g^{\prime}(u^{k-1}_{h}),\chi_{h})\\ =(\tilde{m}(c^{k-1}_{h})h^{\prime}(u^{k-1}_{h}),\chi_{h}),\end{split}$ (5.1a)
@@
where $\tilde{m}(c)=C_{1}(\operatorname{e}\!^{-C_{2}}-c\operatorname{e}\!^{C_{2}})$. The physical parameters are defined, together with their values in the computations, in Table 1. We note that (5.1a)–(5.1c) are linear in $u^{k}$, $\phi^{k}$ and $c^{k}$ respectively. Also, the term $\sigma^{\prime}(u^{k}_{h})\partial_{x_{1}}u^{k}_{h}$ is removed from (5.1b) and incorporated in the boundary conditions for $\phi^{k}_{h}$.
@@ 5.2. Numerical examples
$\hskip 85.35826pt-(\beta u_{h}^{k}(1-u^{k}_{h})X,\chi_{h}),\penalty 10000\ \penalty 10000\ \forall\penalty 10000\ \chi_{h}\in V_{h},$ (5.7)
@@
Electrochemical reaction rate coefficients $C_{1}$ & $C_{2}$ $1/15$ & $1.9345$